Summary of qsymbols

Kristoffer H. Rose^{∗}

Version 1.5 ⟨1994/10/28⟩

Abstract qsymbols is a L^{A}TEX [1] package defining systematic mnemonic abbreviations, starting with a single open quote, `, for symbols and arrows from the amssymb and stmaryrd packages. Optionally a very large class of arrows can be typeset using XY-pic [2].

Contents

1 Introduction 1

2 Simple symbols 2 2.1 Greek letters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Common symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.3 Circled and Boxed Symbols . . . . . . . . . . . . . . . . . . . . . 2 2.4 Bold symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.5 Adding new symbols . . . . . . . . . . . . . . . . . . . . . . . . . 2

3 Orderings 3

4 Arrows 3 4.1 Canned arrows . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 4.2 Adding new arrows: . . . . . . . . . . . . . . . . . . . . . . . . . 3 4.3 Using XY-pic for arrows: . . . . . . . . . . . . . . . . . . . . . . . 3

1 Introduction

qsymbols sets up a number of mnemonic and compact abbreviations for math symbols from L^{A}TEX and the packages amssymb and stmaryrd, which it loads. The abbreviations all start with the backquote character `. Some are a single character, some a more complicated pattern, but always the idea is to use abbreviations that hint at the visual appearance of the symbol. Finally it is possible for the user to add more abbreviations of the simpler categories.

You can retrieve qsymbols as well as the amssymb and stmaryrd packages by anonymous ftp from all CTAN archives in directory /tex-archive/macros/ latex/contrib/supported/^{1} (each package has its own subdirectory).

The following sections describe the various classes of symbols that are loaded.

∗E-mail: krisdiku.dk, W3 URL: http://www.diku.dk/users/kris/.

1The ‘home’ of qsymbols is ftp.diku.dk in directory /diku/users/kris/TeX/.

2 Simple symbols

2.1 Greek letters

All the standard greek letters used in math are available as ` followed by a letter: either lowercase:

          x a b c d e f g h i j k l m n p q r s t w x y z

`x α β χ δ ϵ ϕ γ η ι ψ κ λ μ ν π θ ρ σ τ ω ξ υ ζ

or uppercase:

               x D  F  G  J  L  P  Q  S  W  X  Y

`x Δ Φ Γ Ψ Λ Π Θ Σ Ω Ξ Υ

2.2 Common symbols

Simple symbols are available using ` followed by a symbolic representation of the symbol. The most common have single character representations:

\begin{aligned}x & + & * & : & ; & / & U & - & _ & o & O & . & = & ~ & E & A & ! & ^ & V \\ `x & ± & × & ∈ & ∉ & \ & ∪ & ⊢ & ⊥ & ∘ & ∅ & · & ≡ & ∼ & ∃ & ∀ & ¬ & ∧ & ∨ \end{aligned}

2.3 Circled and Boxed Symbols

These are represented using round and square brackets, respectively:

       x  +  - `* / | `/ `. * `o `^`V <  > ? ! :- a

`(x) ◯ ⊕ ⊖ ⊗ ⊘ ⦶ ⍉ ⊙ ⊛ ⊚  ⧀ ⧁ ∘ • ◯_{¨⌣} ◯^{a}
`[x] □ ⊞ ⊟ ⊠ ⧄ ◫ ⧅ ⊡ ⧆ ⧇ □^{∧} □^{∨} □^{<} □^{>}  ◯_{¨⌢} □^{a}

As it can be seen, ‘undefined’ codes like `(a) and `[a] result in the contents being circled and boxed, respectively.

2.4 Bold symbols

The AMS-L^{A}TEX \boldsymbol command is available by using the special abbreviation `"x for the bold version x of x as well as `"`x where x is on one of the forms described in this section, i.e., `"`a gives α.

2.5 Adding new symbols

Symbols of all the above forms can be added using the form

                  \newqsymbol {`code} {expansion}

which makes `code behave as expansion in math mode. code should be either a single character or some characters in (), [], or {}.

3 Orderings

Two to four consecutive `s indicate an ordering relation.

   ϵ, ϶ ``ϵ ``/ϵ ``ϵ= ``/ϵ= ```ϵ ````ϵ ``϶ ``/϶ ``϶= ``/϶= ```϶ ````϶

<, > < ≮ ≤ ≰ ∧ ⋀  > ≯ ≥ ≱ ∨ ⋁
(,) ⊂ ⊄ ⊆ ⊈ ∩ ⋂  ⊃ ⊅ ⊇ ⊉ ∪ ⋃
[,] ⊏ ⊏̸ ⊑ ⋢ ⊓ ⨅  ⊐ ⊐̸ ⊒ ⋣ ⊔ ⨆
\{, \} ≺ ⊀ ⪯ ⪯̸ ⋏ ⋏  ≻ ⊁ ⪰ ⪰̸ ⋎ ⋎
\<, \> ◁ ⋪ ⊴ ⋬ △ △  ▷ ⋫ ⊵ ⋭ ▽ ▽
~, \~ ∼ ≁ ≃ ≄ ≀ ∽ ∽̸ ⋍ ⋍̸ ∫ ∫
-, \- ⊢ ⊬ ⊧ |≠ ⊤  ⊣ ⊣̸ =| ≠| ⊥

Some abbreviations are provided for convenience:

                      \begin{matrix}x & U & ^ & V & S & P \\ ``x & ⋃ & ⋀ & ⋁ & ∑ & ∏ \end{matrix}

There is no simple way to add more orderings.

4 Arrows

Double quotes "… " make it possible to typeset arrows.

4.1 Canned arrows

The available arrows are shown in figure 1. A notation similar to the XY-pic [2] ‘arrow’ feature notation variant-tail-shaft-tip is used to.

4.2 Adding new arrows:

You can add more ‘canned’ arrows of this kind with commands

                 \newqsymbol {"arrow"} {expansion}

which makes "arrow" behave as expansion in math mode.

4.3 Using XY-pic for arrows:

If the option [xy] is used in the \usepackage command, or if XY-pic [2]^{2} is already loaded, then the XY-pic arrow feature (with the ‘cmtip’ extension) is used to allow a much more general class of arrows, constructed according to the following rules (with examples where they extend the notation used for the standard):

• Basic arrows are composed by combining the variants 23^_, the tips <>|xo`', and the connectors -=.:~.

2XY-pic version 2.12 or later is needed for this to work.

 "<-" ← "<-!" ⟵ "->" → "-!>" ⟶
 "<=" ⇐ "<=!" ⟸ "=>" ⇒ "=!>" ⟹
 "<3" ⇚  "3>" ⇛
[figure: "</-"; "-/>"]
"</=" ⇍  "=/>" ⇏
[figure: "<->"; ↔; "<-!>"; ⟷; "</->"]
"<=>" ⇔ "<=!>" ⟺ "</=>" ⇎
"<-|" ↤ "<-!|" ⟵ "|->" ↦ "|-!>" ⟼
"<=|" ⤆ "<=!|" ⟸ "|=>" ⤇ "|=!>" ⟹
"<-'" ↩  "`->" ↪
"^<-" ↼ "_<-" ↽ "^->" ⇀ "_->" ⇁
"<--" ←−−  "-->" −−→
"<<-" ← ←  "->>" → →
"<<=" ⇐ ⇐  "=>>" ⇒ ⇒
             [figure: "<~>"; "~>"]
[figure: "<-<"; ">->"]
                           "-o" ⊸
                          "->o" → ⊸
                          "->>o" → → ⊸
"<|-" ⇽ "<|-|>" ⇿ "-|>" ⇾

Figure 1: Standard arrow symbols

• The character / ‘negates’ the arrow (once or twice) similar to the way \not does for relations, and the characters ?+* are used to mark the arrow with a symbol corresponding to the reflexive, transitive, and reflexive transitive, closures:

           \begin{matrix}x & `-/> & -//> & =>? & `->+ & |-->* \\ "x" & [figure] & [figure] & [figure: ϵ] & [figure: +] & [figure: ∗] \end{matrix}

• Each ! character makes the arrow a bit longer.

               \begin{matrix}x & -> & -!> & -!!> & & -!!!> \\ "x" & → & ⟶ & & [figure] & & [figure] \end{matrix}

Note: Some arrows are automatically made a bit longer, e.g., the <~> arrow shown above.

• The form {ℓ} adds the ℓ material to the end of the XY-pic arrow. All XY-pic ⟨labels⟩ can be used as described in [2, §16], for example,

        \begin{matrix}x & -\{_1\}> & =\{^\sin\}!>> & =\{|\{`b\}\}!!|> \\ "x" & [figure: 1] & [figure: sin] & [figure: β] \end{matrix}

Use this with care!

• The forms (x) and [x] insert a break with x in a circle/box, respectively:

                   \begin{matrix}x & (1)> & [1]>> \\ "x" & [figure: ◯^{1}] & [figure: □^{1}] \end{matrix}

References

[1] Leslie Lamport. L^{A}TEX—A Document Preparation System. Addison-Wesley, 2nd edition, 1994.

[2] Kristoffer H. Rose and Ross Moore. XY-pic reference manual. Mathematics Report 94–155, MPCE, Macquarie University, NSW 2109, Australia, June 1994. For version 2.10+. Latest version available by anonymous ftp in

ftp.diku.dk:  /diku/users/kris/TeX/xyrefer.ps.Z.